OpenStudy (annie96):

2. Consider the function f(x)=1/4 x+2. a. Find the inverse of f(x) and name it g(x). Show and explain your work. b. Use function composition to show that f(x) and g(x) are inverses of each other.

OpenStudy (anonymous):

Have you tried it yourself?

OpenStudy (annie96):

Yes

OpenStudy (anonymous):

Do you have real picture of it or it it on paper?

OpenStudy (annie96):

it doesn't have a picture it's on paper.

OpenStudy (anonymous):

Is this a inverse function for f(x) and g(x). Can I have a picture by possible?

OpenStudy (anonymous):

I'll just start doing it then.

OpenStudy (fibonaccichick666):

can you show us your work to begin please?

OpenStudy (annie96):

f^−1(x)=4x−8

OpenStudy (fibonaccichick666):

I agree, so that is our g(x)

OpenStudy (fibonaccichick666):

so, now, I bet austin can tell you what function composition is, let's allow him part b.

OpenStudy (annie96):

@FibonacciChick666 do I put g instead of f?

OpenStudy (anonymous):

Replace f(x) with y. y=\[y=\frac{ 1 }{ 4}\times x + 2\] Interchange the variables. \[x=\frac{ 1 }{ 4 }\times y + 2\] Since y is on the right-hand side of the equation, siwtch the sides so it is on the left hand side of the equation. \[\frac{ 1 }{ 4}\times y + 2 =x\] Multiply\[\frac{ 1 }{ 4 }\] by Y to get \[(\frac{ 1 }{ 4 })y\] \[(\frac{ 1 }{ 4 })y+2=x\] Simplify \[(\frac{ 1 }{ 4 })y\] \[\frac{ y }{ 4 }+2=x\] Since 2 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 2 from both sides. \[\frac{ y }{ 4 }=-2+x\] Multiply both sides of the equation by 4. \[y=-2\times(4)+x \times(4)\] Multiply -2 by 4 to get -8. \[y=-8+x \times(4)\] Multiply x by 4 to get x(4). \[y=-8+x(4)\] Multiply x by 4 to get \[x \times 4\] \[y=-8+x \times 4\] Move 4 to the left of the expression \[x \times 4\] \[y=-8+4\times x\] Multiply 4 by x to get 4x \[y=-8 + 4x\] Reorder -8 and 4x. \[y=4x-8\] \[y=4x-8\] @Annie96

OpenStudy (anonymous):

Do you understand it so far

OpenStudy (annie96):

Yes Thank You.

OpenStudy (anonymous):

Do you need more

OpenStudy (anonymous):

Hold on I need to show you more that might change your answer give me like 5 minutes plz

OpenStudy (annie96):

okay

OpenStudy (fibonaccichick666):

oh, yea, you are just renaming f^-1 as g since it's easier to write

OpenStudy (annie96):

okay

OpenStudy (fibonaccichick666):

it's like giving it a nickname

OpenStudy (fibonaccichick666):

now, do you know how to compose functions? Do you know what that means?

OpenStudy (anonymous):

Replace y with \[f-1(x).\] \[f-1(x)=4x-8\] Setup the composite result function. The f o g notation is interpreted as f(g(x)). \[f(g(x))\] Evalute f(g(x)) by substituting in the value of g into f. \[f(4x-8)=\frac{ 1 }{ 4 }\times (4x-8) + 2\] Multiply 1/4 by 4x-8 to get (1/4)(4x-8). \[f(4x-8)=(\frac{ 1 }{ 4 })(4x-8)+2\] Apply the distributive property. \[f(4x-8)=(\frac{ 1 }{ 1 })(\frac{ x }{ 1 })+(\frac{ 1 }{ 4 })(-8)+2\] Simplify. \[f(4x-8)=x+(\frac{ 1 }{ 4 })(-8)+2\] Simplify(1/4)(-8) \[f(4x-8)=x+\frac{ -8 }{ 4}+2\] Divide -8 by 4 to get -2. \[f(4x-8)=x-2+2\] Add-2 and 2 to get 0. \[f(4x-8)=x+0\] Add x and 0 to get x \[f(4x-8)=x\] Since f(g(x)) - x,f-1(x)=4x-8 is the inverse of f(x) =1/4x X+2 \[f-1(x)=4x-8\] \[g-1(x)=4x-8\] I renamed f(x) as g(x) so this should help your answers.

OpenStudy (anonymous):

@Annie96

OpenStudy (anonymous):

I did detailed instructions with inverse functions for f(x) and g(x) so you should understand this and figure this out.

OpenStudy (annie96):

thank you

OpenStudy (anonymous):

I can help you figure out more questions like this or depends on the questions.

OpenStudy (anonymous):

A. I did tons of work and details for f(x) and g(x). B. Gave you the f(x) and g(x) functions

OpenStudy (annie96):

thank you